No relevant resource is found in the selected language.

This site uses cookies. By continuing to browse the site you are agreeing to our use of cookies. Read our privacy policy>

Reminder

To have a better experience, please upgrade your IE browser.

upgrade

In the Case of Fiber Cut, Optical Power Is Normal but SL16 Board Generates B1_SD and B2_SD Alarms

Publication Date:  2012-07-24 Views:  52 Downloads:  0
Issue Description

 The client replaces the fiber, checks the optical power and finds that the optical power is within the normal range but the SL16 board reports the B1_1SD and B2_SD alarms. After all the relevant optical ports are cleaned, the alarms persist. 

 

Alarm Information

Null

Handling Process

1.Replace the original fiber back. It is found that the optical power changes slightly and the bit errors disappear.
2.Check the models of the optical interface board and the fiber. It is found that the optical interface board operates on the 1310 nm wavelength band. Before the cutover, the G.652 fiber is used whereas after the cutover, the G.655 fiber is used. The G.652 fiber can transfer optical signals on the 1310 nm and 1550 nm wavelength bands whereas the G.655 fiber can transfer only optical signals on the 1550 nm wavelength band. The 1310 nm wavelength band is beyond the end wavelength of the G.655 fiber. As a result, bit errors occur in the signal.  

Root Cause

 The alarms are generated after the fiber is replaced. Hence, the problem should be not with the optical interface board but with the fiber. 
 

Suggestions

Normally, this problem does not occur when the G.655 is used to transfer an optical signal on the 1310 wavelength band. In terms of the technical specifications, however, the optical signal on the 1310 nm wavelength band is beyond the end wavelength of the G.655 fiber. As a result, the normal transfer of the optical signal is not ensured. Note to prevent this problem when using the G.655 fiber.  

See also case in chinese: SC0000166736 
 

END